lizzie777:

In a football tournament, the Bees scored 9 less than three times as many points as the Hornets. The Wasps scored 28 more points than the Hornets. Together the three teams scored 184 points. Let x represent the number of points scored by the Hornets. Complete the steps to solve the equation (3x – 9) + x + (x + 28) = 184. Then answer the question. 1. Combine like terms. 2. Use the subtraction property of equality. 3. Use the division property of equality. 4. Use the variable to answer the question. How many points did the Bees score? points

darkknight:

Do you know how to combine like terms?

lizzie777:

yea

lizzie777:

but not really good at it

darkknight:

So we can combine the x terms with each other and the non variable terms with each other. (3x – 9) + x + (x + 28) = 184 So we have 3x+x+x-9+28 = 184

darkknight:

Can you combine these? 3x+x+x-9+28 = 184?

lizzie777:

yea

darkknight:

So now we have 5x+19=184

darkknight:

Do you know what the subtraction property of equality is?

lizzie777:

a liitle bit

lizzie777:

I'm not that good in math

darkknight:

So its states that if we subtract from one side of an equation, we also must subtract the same value from the other side of the equation to keep the equation the same.

darkknight:

The goal here is to isolate x 5x+19=184 We subtract 19 from both sides 5x+19-19 = 184-19 5x = 165

darkknight:

Now do you know what the division property of equality is?

lizzie777:

yea

lizzie777:

isn't it where you divide on both sides

darkknight:

Yes it is

darkknight:

So we want to divide by 5 on both sides So 5x/5= 165/5 x = 31

darkknight:

Now can you answer the last part by yourself?

lizzie777:

thank you

darkknight:

You still have the last part to answer. xD

darkknight:

(3x – 9) + x + (x + 28) = 184. Is our equation We have to remember that this is the total score of the Bees, hornets and wasps respectively. So 3x-9 is the points that the bees scored. Plug in the value you got for x in there and find the points that they scored

lizzie777:

the answer is 90

darkknight:

Not quite. So we got 31 for x Substituting, we get 3(31)-9 93-9 84

darkknight:

Do you see how I got that?

lizzie777:

I put that and it said it is 90

lizzie777:

yea I put 84 first and it said it was wrong

darkknight:

Okay, i am not very smart. 165/5 = 33. My bad...

darkknight:

I don't know how I messed that up. Apologies, so we plug in 33 instead of 31

lizzie777:

its all good

darkknight:

Now can you solve it? Plug in 33 for x in 3x-9

darkknight:

So yeah, the answer was 90. You were right

darkknight:

Keep it up :)

lizzie777:

thanks

darkknight:

No problem